OpenStudy (anonymous):

Identify all of the following solutions of sqrtx-8 + 8 = x . x = 8 x = 9 x = 8 and x = 9 None of the above

OpenStudy (angela210793):

is it x-8 under sqr or just x?

OpenStudy (anonymous):

\[\sqrt{x-8}+8 = x\] \[\sqrt{x-8} = x-8\] Let k = x-8. \[\sqrt{k} = k\] \[k = 0,1\] Thus \[x-8= 0,1\] \[x= 8,9\]